Use the given conditions to write an equation for the line in point-slope form and in slope-intercept form Passing through (-8,-6) and parallel to the line whose equation is y-5x+1 Write an equation for the line in point-slope form. (Simplify your answer. Use integers or fractions for any numbers in the equation.) Write an equation for the line in slope-intercept form. (Simplify your answer. Use integers or fractions for any numbers in the equation.) Enter your answer in each of the answer boxes OType here to search
Answer #1

Given that the line passes through the point (-8,-6)

and it is parallel to the line y = -5x+1

we know that two parallel lines will have same slope.

So, the line equation, y = mx+c, we can conclude that m = -5

point slope form

y-y1 = m(x-x1)

y+6 = -5(x+8)

y+6 = -5x-40

Slope intercept form:

y = -5x-46

where slope = -5 and y-intercept = -46
Answer #2

SOLUTION :


Line  to which the required line is parallel is given as : y = - 5x + 1 


Its slope = - 5 


Slope of the required line  = - 5


Since point (- 8, - 6) is on this line,


Equation of the required line in point slope form will be :


(y - (- 6) ) = - 5( x - (- 8) )

=> (y + 6) = - 5(x + 8). ANSWER).


Converting it into slope intercept form :

=> y = - 5x - 40 - 6 

=> y = - 5x - 46


So, equation of the required line in slope-intercept form :


y = - 5x - 46   (ANSWER).

answered by: Tulsiram Garg